# Abelian surfaces over finite fields containing no curves of genus $3$ or less

### v3

@@ 2. Abelian surfaces containing genus $3$ curves
    $A$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$.
@@
[…]
@@
If $A$ does not contain an absolutely irreducible curve of geometric genus $2$, then, from Lemma 2.1, we deduce that the genus $3$ $\mathbb{F}_{q}$-irreducible curve in the statement of Theorem 2.2 is either an absolutely irreducible smooth curve of genus $3$ or an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$ which is not absolutely irreducible.
@@
If an abelian surface $A$ in $\mathcal{C}$ contains an absolutely irreducible curve $C$ of geometric genus $3$, then $A$ is an isogeny factor of the Jacobian $\mathop{\rm Jac}\nolimits(\tilde{C})$ of the normalisation $\tilde{C}$ of $C$ by [9, Prop. 2] [^3] The author requires the curve to have a rational point on the smooth curve to guarantee the existence of an embedding into its Jacobian. However, the existence of a divisor of degree $1$ is sufficient. A curve defined over a finite field always admits such a divisor.. Hence, the Weil polynomial of $\mathop{\rm Jac}\nolimits(\tilde{C})$ must be of the form $(t^{2}-2)^{2}f_{E}(t)$, where $f_{E}(t)$ is the Weil polynomial of an elliptic curve over $\mathbb{F}_{2}$.
@@ 3. Factorisation of $2$
  • If $b=1-2q$ then $12q-b=4q-$ 1. So $c$ is odd. Hence $d\equiv 3\bmod 4$.
  • If $b=2-2q$ then $22q-b=4q-$ 2. So $c$ is odd. Hence $d\equiv 2\bmod 4$.
  • If $b=-q$ and $q$ is a square then $d=3$.
  • If $b=-q$, $p=2$ and $q$ is not a square then $d=6\equiv 2\bmod 4$.
@@
Note that $R=R^{+}[\pi]$ (resp. $\mathcal{O}^{+}[\pi]$) is locally free of rank $2$ over $R^{+}$ (resp. $\mathcal{O}^{+}$). Hence $[\mathcal{O}^{+}:R^{+}]^{2}=[\mathcal{O}^{+}[\pi]:R]$ divides $[\mathcal{O}:R]$. Therefore,
@@
Assume first that $f$ is non-ordinary. Then $f(t)=t^{4}-qt^{2}+q^{2}$ by Theorem 1.4 and Lemma 1.12. Note that $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $\mathbb{Q}(\zeta_{3})$ in $K$ (cf. [15, Sec. 4]). Moreover, as we have shown in the proof of Lemma 3.1, we have that $K^{+}=\mathbb{Q}(\sqrt{3})$ or $K^{+}=\mathbb{Q}(\sqrt{6})$. In the first case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{3})$ and one computes that $\Delta_{K^{+}}=12$ and $\Delta_{K}=12^{2}$. In the second case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{6})$ and one computes that $\Delta_{K^{+}}=24$ and $\Delta_{K}=24^{2}$. In both cases $K/K^{+}$ is unramified.
Hence, $f(t)$ must be ordinary. Then, by Lemma 1.12, either $b=1-2q$ or, $b=2-2q$ and $q$ is odd. In the first case, Equation (3) implies that $N=1$, which cannot happen by assumption. So we must be in the second case, that is, iv holds. Equation (3) implies that if $b=2-2q$ (with $q$ odd) then $N\in\{1,2,4\}$, and in fact $N\in\{2,4\}$ since we are assuming that $K/K^{+}$ is ramified. Since $\mathfrak{m}$ is the unique maximal ideal of $\mathcal{O}^{+}$ above $2$ by Lemma 3.1, we obtain that $\mathfrak{m}$ ramifies in $K$, that is, ii holds.
@@
If $K/K^{+}$ is ramified or if there exists a prime in $K^{+}$ that divides $(\pi-q/\pi)$ and which stays inert in $K$, then, by [14, Thm. 1.1], there is a principally polarised abelian surface in the isogeny class associated to $f(t)$. Hence $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ by Theorem 1.4.
@@
Assume by contradiction that $f(t)$ is ordinary. Then, by Lemma 1.12, either $b=1-2q$, or $b=2-2q$ with $q$ odd. If $b=1-2q$ then $(\pi-q/\pi)^{2}=-1$ which implies that $1\in\mathfrak{n}^{2}\mathcal{O}$. This is not possible since $\mathfrak{n}\mathcal{O}$ is a maximal ideal of $\mathcal{O}$ by assumption. If $b=2-2q$ with $q$ odd, then $2\in\mathfrak{n}^{2}\mathcal{O}$. By Lemma 3.1, it follows that $\mathfrak{n}$ is the unique maximal ideal of $\mathcal{O}^{+}$ above $2$. By Lemma 3.2, $\mathfrak{n}$ is ramified in $K$, leading to a contradiction also in this case. Therefore $f(t)$ is non-ordinary. ∎
@@
If $f(t)$ is ordinary then [13, proof of Lemma 12.2] shows that $K$ contains $M=\mathbb{Q}(\zeta_{3})$. If $f(t)$ is not ordinary, then [15, Thm. 2] tells us that
@@
Then $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $M$ in $K$ (cf. [15, Sec. 4]). ∎
@@ 4. Kernels of polarisations of abelian varieties
Let $G(\mathop{\rm Mod}\nolimits_{S})$ be the Grothendieck group of $\mathop{\rm Mod}\nolimits_{S}$ which is defined as the quotient of the free abelian group on the isomorphism classes of objects in $\mathop{\rm Mod}\nolimits_{S}$ by the subgroup generated by the expressions $M-M^{\prime}-M^{\prime\prime}$ for all exact sequences $0\to M^{\prime}\to M\to M^{\prime\prime}\to 0$ in $\mathop{\rm Mod}\nolimits_{S}$. For $M\in\mathop{\rm Mod}\nolimits_{S}$, we denote its class in $G(\mathop{\rm Mod}\nolimits_{S})$ by $[M]_{S}$. The association $M\mapsto|M|$ induces a group homomorphism $G(\mathop{\rm Mod}\nolimits_{S})\to\mathbb{Q}^{*}$. Note that $G(\mathop{\rm Mod}\nolimits_{S})$ is a free abelian group on the simple objects of $\mathop{\rm Mod}\nolimits_{S}$. Every such simple object is $S$-linearly isomorphic to $S/\mathfrak{N}$ for a maximal $S$-ideal $\mathfrak{N}$. An element of $G(\mathop{\rm Mod}\nolimits_{S})$ is called effective if it is a sum of positive multiples of simple objects. Let $\alpha=a/b$ be an element of $K^{\times}$. The principal element generated by $\alpha$ is the element $\mathop{\rm Pr}\nolimits_{S}(\alpha)=[S/aS]_{S}-[S/bS]_{R}$ of $G(\mathop{\rm Mod}\nolimits_{S})$. Note that $\mathop{\rm Pr}\nolimits_{S}$ induces a group homomorphism from $K^{\times}$ to $G(\mathop{\rm Mod}\nolimits_{S})$.
@@
Let $G(\mathop{\rm Ker_{\mathcal{C}}}\nolimits)$ be the associated Grothendieck group, which is defined analogously to the one of an order. For an element $G$ in $\mathop{\rm Ker_{\mathcal{C}}}\nolimits$ we denote its class in $G(\mathop{\rm Ker_{\mathcal{C}}}\nolimits)$ by $[G]_{\mathcal{C}}$. Moreover, we denote by $\widehat{G}$ its Cartier dual.
@@
  • set $\epsilon([\alpha_{p}]_{\mathcal{C}})=[M]_{R}$ where $M=\mathbb{Z}/p\mathbb{Z}$ with the $R$-module structure induced by letting $\pi$ and $q/\pi$ act as 00.
@@
Let $\varphi:A\to B$ be an isogeny between elements of $\mathcal{C}$. Let $M$ be a finite $R$-module such that $\epsilon([\ker\phi]_{\mathcal{C}})=[M]_{R}$. Then
@@
$[G]_{\mathcal{C}}=[\alpha_{p}^{n}\times G^{\prime}]_{\mathcal{C}}=n[\alpha_{p}]_{\mathcal{C}}\oplus[G^{\prime}]_{\mathcal{C}}.$
@@
Let $S$ be an order in $K$ such that $S=\overline{S}$ and that $S$ is locally free of rank $2$ over $S^{+}=S\cap K^{+}$. Then $Z(S)=t_{S/S^{+}}(G(\mathop{\rm Mod}\nolimits_{S^{+}}))$. Moreover, if $M\in\mathop{\rm Mod}\nolimits_{S}$ is such that $[M]_{S}\in Z(S)$ then $[M]_{S}=t_{S/S^{+}}(P^{\prime})$ for some effective $P^{\prime}\in G(\mathop{\rm Mod}\nolimits_{S^{+}})$.
@@
because $S$ is locally free of rank $2$ over $S^{+}$. Hence $N_{S/S^{+}}([S/\mathfrak{n}S]_{S})\in 2G(\mathop{\rm Mod}\nolimits_{S^{+}})$. Since $S=\overline{S}$ and $\mathfrak{n}S=\overline{\mathfrak{n}S}$, by Lemma 4.1, we get that $[S^{+}/\mathfrak{n}\otimes_{S^{+}}S]_{S}=[S/\mathfrak{n}S]_{S}$ is symmetric. This concludes the proof that
@@
$[S/\mathfrak{N}\oplus S/\overline{\mathfrak{N}}]_{S}=[S/\mathfrak{N}\overline{\mathfrak{N}}]_{S}=[S/\mathfrak{n}S]_{S}=t_{S/S^{+}}([S^{+}/\mathfrak{n}]_{S^{+}}).$
@@
$[S/\mathfrak{N}]_{S}=[S/\mathfrak{n}S]_{S}=t_{S/S^{+}}([S^{+}/\mathfrak{n}]_{S^{+}}).$
@@
    There exists a finite $R$-module $M$ of order $|M|=\ell^{2}$ such that $[M]_{R}$ is in $Z(R)$.
@@
If any of the equivalent statements hold then $[M]_{R}=t_{R/R^{+}}([R^{+}/\mathfrak{l}]_{R^{+}})$ where $\mathfrak{l}$ is the unique maximal ideal of $R^{+}$ above $\ell$ corresponding to any linear factor of $f^{+}(t)\bmod\ell$.
@@
This is part of [14, Prop. 6.2]. ∎
@@
The group $H(R)$ is a vector space over $\mathbb{F}_{2}$ whose basis is given by the classes of the simple $R$-modules $R/\mathfrak{n}$ where $\mathfrak{n}$ is a maximal $R$-ideal such that $\mathfrak{n}=\overline{\mathfrak{n}}$ and the index $[R/\mathfrak{n}:R^{+}/(R^{+}\cap\mathfrak{n})]$ of residue fields is even. Such maximal ideals are called the generating primes of $H(R)$.
@@
This is a special case of [14, Prop. 6.4]. ∎
@@
This is part of [14, Prop. 7.1]. ∎
@@ 5. Polarisations of degree $4$ on abelian surfaces
$[M]_{R}=[R^{+}/\mathfrak{l}\otimes_{R^{+}}R]_{R}=[R/\mathfrak{l}R]_{R},$
where $\mathfrak{l}$ is a maximal ideal of $R^{+}$ such that $R^{+}/\mathfrak{l}\simeq\mathbb{F}_{2}$. By Lemma 3.3, $K/K^{+}$ is unramified. So, the Artin map induces an isomorphism $\mathop{\rm Gal}\nolimits(K/K^{+})\simeq\mathcal{B}(\mathcal{O})$ by Proposition 4.8. Let $\alpha$ be the non-zero element of $\mathcal{B}(\mathcal{O})$, which corresponds to the Artin symbol of any maximal ideal of $\mathcal{O}^{+}$ that stays inert in $K$. By Theorem 4.10 and Proposition 4.11, $I_{\mathcal{C}}$ is non-zero and in $i^{*}(\mathcal{B}(\mathcal{O}))$. Hence $I_{\mathcal{C}}=i^{*}(\alpha)$. Let $z_{\mathfrak{l}}$ be the image of $[R/\mathfrak{l}R]_{R}$ in $\mathcal{B}(R)$. Note that
@@
where $x_{\mathfrak{l}}$ is the image of $[R/\mathfrak{l}R]_{R}$ in $H(R)$. We obtain that $z_{\mathfrak{l}}-I_{\mathcal{C}}=(\Psi\oplus i^{*})((x_{\mathfrak{l}},\alpha))$. Hence there is an $R$-module $M$ of order $4$ which is attainable if and only if
@@
Assume now that $2$ is not inert in $K^{+}$. Let $\mathfrak{M}$ be a maximal ideal of $\mathcal{O}$ containing $\mathfrak{l}R$. Denote the image of $[\mathcal{O}/\mathfrak{M}]_{\mathcal{O}}$ in $H(\mathcal{O})$ by $y_{\mathfrak{M}}$. By Proposition 3.5, we have that $\mathfrak{M}=\mathfrak{m}\mathcal{O}$ for a maximal ideal $\mathfrak{m}$ of $\mathcal{O}^{+}$. This implies that $\chi(y_{\mathfrak{M}})=\alpha$. So, to prove that $R/\mathfrak{l}R$ is attainable, we are left to show that $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. We first consider the case that $\mathfrak{l}R$ is a maximal ideal of $R$. Note that $\mathfrak{l}R=\mathfrak{M}\cap R$ and we have an isomorphism $R/\mathfrak{l}R\simeq\mathcal{O}/\mathfrak{M}$. Hence $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. Therefore, $R/\mathfrak{l}R$ is attainable. Now we consider the case when $\mathfrak{l}R$ is not a maximal ideal of $R$. Set $\mathfrak{L}=\mathfrak{M}\cap R$. The natural surjective map $R/\mathfrak{l}R\twoheadrightarrow R/\mathfrak{L}$ is not an isomorphism. Since $\mathfrak{L}=\overline{\mathfrak{L}}$ and $[R/\mathfrak{l}R]_{R}\in Z(R)$ we get that
$[R/\mathfrak{l}R]_{R}=2[R/\mathfrak{L}]_{R}.$
Moreover, also $[\mathcal{O}/\mathfrak{M}]_{R}=2[R/\mathfrak{L}]_{R}$. Hence, again, $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. Therefore, as before, $R/\mathfrak{l}R$ is attainable. ∎
@@
$[M]_{R}=[R^{+}/\mathfrak{l}\otimes_{R^{+}}R]_{R}=[R/\mathfrak{l}R]_{R}.$
Since $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, the obstruction element $I_{\mathcal{C}}$ is trivial by Theorem 4.10. Hence, there is an $R$-module $M$ of order $4$ which is attainable if and only if the image $z_{\mathfrak{l}}$ of $[R/\mathfrak{l}R]_{R}$ in $\mathcal{B}(R)$ is trivial. Note that
@@
where $x_{\mathfrak{l}}$ is the image of $[R/\mathfrak{l}R]_{R}$ in $H(R)$. Also, the $R$-module $R/\mathfrak{l}R$ has length at most $2$ and that it has length $1$ if and only if $\mathfrak{l}R$ is a maximal ideal of $R$.
If the length of $R/\mathfrak{l}R$ is $2$ then $[R/\mathfrak{l}R]_{R}=[R/\mathfrak{n}]_{R}+[R/\overline{\mathfrak{n}}]_{R}$ for a maximal ideal $\mathfrak{n}$ of $R$, with possibly $\mathfrak{n}=\overline{\mathfrak{n}}$, and so, by Lemma 4.1, $[R/\mathfrak{l}R]_{R}$ is in $B(R)$. Hence, $x_{\mathfrak{l}}$ is trivial in $H(R)$ and so $z_{\mathfrak{l}}=I_{\mathcal{C}}$ in $\mathcal{B}(R)$. This means that $R/\mathfrak{l}R$ is an attainable $R$-module of order $4$.

### v4

@@ 2. Abelian surfaces containing genus $3$ curves
    AA contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$.
@@
To prove $i\Rightarrowi$, let $C$ be an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$ lying over $A$. As in the proof of Proposition 1.3, the adjunction formula, Lemma 1.2 and the Nakai–Moishezon criterion imply that $C$ is an ample divisor, which in turn defines a polarisation of degree $4$ by Equation (1). ∎
@@
If $A$ does not contain an absolutely irreducible curve of geometric genus $2$, then, from Lemma 2.1, we deduce that the genus 33 $\mathbb{F}_{q}$-irreducible curve in the statement of Theorem 2.2 is either an absolutely irreducible smooth curve of genus $3$ or an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$ which is not absolutely irreducible.
@@
[…]
@@ 3. Factorisation of 22
  • If b=12qb=1-2q then 2qb=4q12q-b=4q-1. So $c$ is odd. Hence $d\equiv 3\bmod 4$.
  • If b=22qb=2-2q then 2qb=4q22q-b=4q-2. So $c$ is odd. Hence $d\equiv 2\bmod 4$.
  • If b=qb=-q and $q$ is a square then $d=3$.
  • If b=qb=-q, $p=2$ and $q$ is not a square then $d=6\equiv 2\bmod 4$.
@@
Note that $R=R^{+}[\pi]$ (resp. $\mathcal{O}^{+}[\pi]$) is locally free of rank $2$ over $R^{+}$ (resp. $\mathcal{O}^{+}$). Hence $\mathcal{O}^{+}:R^{+}]^{2}=[\mathcal{O}^{+}[\pi]:R]$ divides $\mathcal{O}:R]$. Therefore,
@@
Assume first that $f$ is non-ordinary. Then $f(t)=t^{4}-qt^{2}+q^{2}$ by Theorem 1.4 and Lemma 1.12. Note that $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $\mathbb{Q}(\zeta_{3})$ in $K$ (cf. [12, Sec. 4]). Moreover, as we have shown in the proof of Lemma 3.1, we have that $K^{+}=\mathbb{Q}(\sqrt{3})$ or $K^{+}=\mathbb{Q}(\sqrt{6})$. In the first case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{3})$ and one computes that $\Delta_{K^{+}}=12$ and $\Delta_{K}=12^{2}$. In the second case, we have $K=\mathbb{Q}(\zeta_{3},\sqrt{6})$ and one computes that $\Delta_{K^{+}}=24$ and $\Delta_{K}=24^{2}$. In both cases $K/K^{+}$ is unramified.
Hence, $f(t)$ must be ordinary. Then, by Lemma 1.12, either b=12qb=1-2q or, b=22qb=2-2q and $q$ is odd. In the first case, Equation (3) implies that $N=1$, which cannot happen by assumption. So we must be in the second case, that is, iv holds. Equation (3) implies that if b=22qb=2-2q (with $q$ odd) then $N\in\{1,2,4\}$, and in fact $N\in\{2,4\}$ since we are assuming that $K/K^{+}$ is ramified. Since $\mathfrak{m}$ is the unique maximal ideal of $\mathcal{O}^{+}$ above $2$ by Lemma 3.1, we obtain that $\mathfrak{m}$ ramifies in $K$, that is, ii holds.
@@
If $K/K^{+}$ is ramified or if there exists a prime in $K^{+}$ that divides $(\pi-q/\pi)$ and which stays inert in $K$, then, by [16, Thm. 1.1], there is a principally polarised abelian surface in the isogeny class associated to $f(t)$. Hence $f\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ by Theorem 1.4.
@@
Assume by contradiction that $f(t)$ is ordinary. Then, by Lemma 1.12, either b=12qb=1-2q, or b=22qb=2-2q with $q$ odd. If b=12qb=1-2q then $(\pi-q/\pi)^{2}=-1$ which implies that $1\in\mathfrak{n}^{2}\mathcal{O}$. This is not possible since $\mathfrak{n}\mathcal{O}$ is a maximal ideal of $\mathcal{O}$ by assumption. If b=22qb=2-2q with $q$ odd, then $2\in\mathfrak{n}^{2}\mathcal{O}$. By Lemma 3.1, it follows that $\mathfrak{n}$ is the unique maximal ideal of $\mathcal{O}^{+}$ above $2$. By Lemma 3.2, $\mathfrak{n}$ is ramified in $K$, leading to a contradiction also in this case. Therefore $f(t)$ is non-ordinary. ∎
@@
If $f(t)$ is ordinary then [15, proof of Lemma 12.2] shows that $K$ contains $M=\mathbb{Q}(\zeta_{3})$. If $f(t)$ is not ordinary, then [12, Thm. 2] tells us that
@@
Then $\zeta_{3}\mapsto-\pi^{2}/q$ gives an embedding of $M$ in $K$ (cf. [12, Sec. 4]). ∎
@@ 4. Kernels of polarisations of abelian varieties
Let $G(\mathop{\rm Mod}\nolimits_{S})$ be the Grothendieck group of $\mathop{\rm Mod}\nolimits_{S}$ which is defined as the quotient of the free abelian group on the isomorphism classes of objects in $\mathop{\rm Mod}\nolimits_{S}$ by the subgroup generated by the expressions $M-M^{\prime}-M^{\prime\prime}$ for all exact sequences $0\to M^{\prime}\to M\to M^{\prime\prime}\to 0$ in $\mathop{\rm Mod}\nolimits_{S}$. For $M\in\mathop{\rm Mod}\nolimits_{S}$, we denote its class in $G(\mathop{\rm Mod}\nolimits_{S})$ by $M]_{S}$. The association $M\mapsto|M|$ induces a group homomorphism $G(\mathop{\rm Mod}\nolimits_{S})\to\mathbb{Q}^{*}$. Note that $G(\mathop{\rm Mod}\nolimits_{S})$ is a free abelian group on the simple objects of $\mathop{\rm Mod}\nolimits_{S}$. Every such simple object is $S$-linearly isomorphic to $S/\mathfrak{N}$ for a maximal $S$-ideal $\mathfrak{N}$. An element of $G(\mathop{\rm Mod}\nolimits_{S})$ is called effective if it is a sum of positive multiples of simple objects. Let $\alpha=a/b$ be an element of $K^{\times}$. The principal element generated by $\alpha$ is the element $\mathop{\rm Pr}\nolimits_{S}(\alpha)=[S/aS]_{S}-[S/bS]_{R}$ of $G(\mathop{\rm Mod}\nolimits_{S})$. Note that $\mathop{\rm Pr}\nolimits_{S}$ induces a group homomorphism from $K^{\times}$ to $G(\mathop{\rm Mod}\nolimits_{S})$.
@@
Let $G(\mathop{\rm Ker_{\mathcal{C}}}\nolimits)$ be the associated Grothendieck group, which is defined analogously to the one of an order. For an element $G$ in $\mathop{\rm Ker_{\mathcal{C}}}\nolimits$ we denote its class in $G(\mathop{\rm Ker_{\mathcal{C}}}\nolimits)$ by $G]_{\mathcal{C}}$. Moreover, we denote by $\widehat{G}$ its Cartier dual.
@@
  • set $\epsilon([\alpha_{p}]_{\mathcal{C}})=[M]_{R}$ where $M=\mathbb{Z}/p\mathbb{Z}$ with the $R$-module structure induced by letting $\pi$ and $q/\pi$ act as 0.
@@
Let $\varphi:A\to B$ be an isogeny between elements of $\mathcal{C}$. Let $M$ be a finite $R$-module such that $\epsilon([\ker\varphi]_{\mathcal{C}})=[M]_{R}$. Then
@@
$G]_{\mathcal{C}}=[\alpha_{p}^{n}\times G^{\prime}]_{\mathcal{C}}=n[\alpha_{p}]_{\mathcal{C}}\oplus[G^{\prime}]_{\mathcal{C}}.$
@@
Let $S$ be an order in $K$ such that $S=\overline{S}$ and that $S$ is locally free of rank $2$ over $S^{+}=S\cap K^{+}$. Then $Z(S)=t_{S/S^{+}}(G(\mathop{\rm Mod}\nolimits_{S^{+}}))$. Moreover, if $M\in\mathop{\rm Mod}\nolimits_{S}$ is such that $M]_{S}\in Z(S)$ then $M]_{S}=t_{S/S^{+}}(P^{\prime})$ for some effective $P^{\prime}\in G(\mathop{\rm Mod}\nolimits_{S^{+}})$.
@@
because $S$ is locally free of rank $2$ over $S^{+}$. Hence $N_{S/S^{+}}([S/\mathfrak{n}S]_{S})\in 2G(\mathop{\rm Mod}\nolimits_{S^{+}})$. Since $S=\overline{S}$ and $\mathfrak{n}S=\overline{\mathfrak{n}S}$, by Lemma 4.1, we get that $S^{+}/\mathfrak{n}\otimes_{S^{+}}S]_{S}=[S/\mathfrak{n}S]_{S}$ is symmetric. This concludes the proof that
@@
$S/\mathfrak{N}\oplus S/\overline{\mathfrak{N}}]_{S}=[S/\mathfrak{N}\overline{\mathfrak{N}}]_{S}=[S/\mathfrak{n}S]_{S}=t_{S/S^{+}}([S^{+}/\mathfrak{n}]_{S^{+}}).$
@@
$S/\mathfrak{N}]_{S}=[S/\mathfrak{n}S]_{S}=t_{S/S^{+}}([S^{+}/\mathfrak{n}]_{S^{+}}).$
@@
    There exists a finite $R$-module $M$ of order $|M|=\ell^{2}$ such that $M]_{R}$ is in $Z(R)$.
@@
If any of the equivalent statements hold then $M]_{R}=t_{R/R^{+}}([R^{+}/\mathfrak{l}]_{R^{+}})$ where $\mathfrak{l}$ is the unique maximal ideal of $R^{+}$ above $\ell$ corresponding to any linear factor of $f^{+}(t)\bmod\ell$.
@@
This is part of [16, Prop. 6.2]. ∎
@@
The group $H(R)$ is a vector space over $\mathbb{F}_{2}$ whose basis is given by the classes of the simple $R$-modules $R/\mathfrak{n}$ where $\mathfrak{n}$ is a maximal $R$-ideal such that $\mathfrak{n}=\overline{\mathfrak{n}}$ and the index $R/\mathfrak{n}:R^{+}/(R^{+}\cap\mathfrak{n})]$ of residue fields is even. Such maximal ideals are called the generating primes of $H(R)$.
@@
This is a special case of [16, Prop. 6.4]. ∎
@@
This is part of [16, Prop. 7.1]. ∎
@@ 5. Polarisations of degree $4$ on abelian surfaces
$M]_{R}=[R^{+}/\mathfrak{l}\otimes_{R^{+}}R]_{R}=[R/\mathfrak{l}R]_{R},$
where $\mathfrak{l}$ is a maximal ideal of $R^{+}$ such that $R^{+}/\mathfrak{l}\simeq\mathbb{F}_{2}$. By Lemma 3.3, $K/K^{+}$ is unramified. So, the Artin map induces an isomorphism $\mathop{\rm Gal}\nolimits(K/K^{+})\simeq\mathcal{B}(\mathcal{O})$ by Proposition 4.8. Let $\alpha$ be the non-zero element of $\mathcal{B}(\mathcal{O})$, which corresponds to the Artin symbol of any maximal ideal of $\mathcal{O}^{+}$ that stays inert in $K$. By Theorem 4.10 and Proposition 4.11, $I_{\mathcal{C}}$ is non-zero and in $i^{*}(\mathcal{B}(\mathcal{O}))$. Hence $I_{\mathcal{C}}=i^{*}(\alpha)$. Let $z_{\mathfrak{l}}$ be the image of $R/\mathfrak{l}R]_{R}$ in $\mathcal{B}(R)$. Note that
@@
where $x_{\mathfrak{l}}$ is the image of $R/\mathfrak{l}R]_{R}$ in $H(R)$. We obtain that $z_{\mathfrak{l}}-I_{\mathcal{C}}=(\Psi\oplus i^{*})((x_{\mathfrak{l}},\alpha))$. Hence there is an $R$-module $M$ of order $4$ which is attainable if and only if
@@
Assume now that $2$ is not inert in $K^{+}$. Let $\mathfrak{M}$ be a maximal ideal of $\mathcal{O}$ containing $\mathfrak{l}R$. Denote the image of $\mathcal{O}/\mathfrak{M}]_{\mathcal{O}}$ in $H(\mathcal{O})$ by $y_{\mathfrak{M}}$. By Proposition 3.5, we have that $\mathfrak{M}=\mathfrak{m}\mathcal{O}$ for a maximal ideal $\mathfrak{m}$ of $\mathcal{O}^{+}$. This implies that $\chi(y_{\mathfrak{M}})=\alpha$. So, to prove that $R/\mathfrak{l}R$ is attainable, we are left to show that $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. We first consider the case that $\mathfrak{l}R$ is a maximal ideal of $R$. Note that $\mathfrak{l}R=\mathfrak{M}\cap R$ and we have an isomorphism $R/\mathfrak{l}R\simeq\mathcal{O}/\mathfrak{M}$. Hence $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. Therefore, $R/\mathfrak{l}R$ is attainable. Now we consider the case when $\mathfrak{l}R$ is not a maximal ideal of $R$. Set $\mathfrak{L}=\mathfrak{M}\cap R$. The natural surjective map $R/\mathfrak{l}R\twoheadrightarrow R/\mathfrak{L}$ is not an isomorphism. Since $\mathfrak{L}=\overline{\mathfrak{L}}$ and $R/\mathfrak{l}R]_{R}\in Z(R)$ we get that
$R/\mathfrak{l}R]_{R}=2[R/\mathfrak{L}]_{R}.$
Moreover, also $\mathcal{O}/\mathfrak{M}]_{R}=2[R/\mathfrak{L}]_{R}$. Hence, again, $N(y_{\mathfrak{M}})=x_{\mathfrak{l}}$. Therefore, as before, $R/\mathfrak{l}R$ is attainable. ∎
@@
$M]_{R}=[R^{+}/\mathfrak{l}\otimes_{R^{+}}R]_{R}=[R/\mathfrak{l}R]_{R}.$
Since $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, the obstruction element $I_{\mathcal{C}}$ is trivial by Theorem 4.10. Hence, there is an $R$-module $M$ of order $4$ which is attainable if and only if the image $z_{\mathfrak{l}}$ of $R/\mathfrak{l}R]_{R}$ in $\mathcal{B}(R)$ is trivial. Note that
@@
where $x_{\mathfrak{l}}$ is the image of $R/\mathfrak{l}R]_{R}$ in $H(R)$. Also, the $R$-module $R/\mathfrak{l}R$ has length at most $2$ and that it has length $1$ if and only if $\mathfrak{l}R$ is a maximal ideal of $R$.
If the length of $R/\mathfrak{l}R$ is $2$ then $R/\mathfrak{l}R]_{R}=[R/\mathfrak{n}]_{R}+[R/\overline{\mathfrak{n}}]_{R}$ for a maximal ideal $\mathfrak{n}$ of $R$, with possibly $\mathfrak{n}=\overline{\mathfrak{n}}$, and so, by Lemma 4.1, $R/\mathfrak{l}R]_{R}$ is in $B(R)$. Hence, $x_{\mathfrak{l}}$ is trivial in $H(R)$ and so $z_{\mathfrak{l}}=I_{\mathcal{C}}$ in $\mathcal{B}(R)$. This means that $R/\mathfrak{l}R$ is an attainable $R$-module of order $4$.